OpenStudy (anonymous):

Meg is 6 years older than Victor. Meg's age is 2 years less than five times Victor's age. The equations below model the relationship between Meg's age (m) and Victor's age (v): m = v + 6 m = 5v − 2 Which is a possible correct method to find Meg's and Victor's ages? a. Solve m + 6 = 5m − 2 to find the value of m. b. Write the points where the graphs of the equations intersect the x axis. c. Solve v + 6 = 5v − 2 to find the value of v. d. Write the points where the graphs of the equations intersect the y axis.

OpenStudy (anonymous):

i had that question!!

OpenStudy (anonymous):

You did?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

sure hold on

OpenStudy (anonymous):

Ok. Thanks so much.

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b

OpenStudy (anonymous):

B is the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you explain why for me? I would like to understand

OpenStudy (anonymous):

hi jhonyy9

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

did you do flvs when you had this question?

OpenStudy (anonymous):

what does fvl means?

OpenStudy (anonymous):

florida virtual school

OpenStudy (anonymous):

yes except that i am in pensylvania

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

and the answer is actually C

jhonyy9 (jhonyy9):

what result you will get for v and m ?

OpenStudy (anonymous):

ok. May you explain please?

OpenStudy (anonymous):

I don't understand the question. That's why I found this website and so yeah.

OpenStudy (anonymous):

cause if we know victors age it would be easy to find megs

jhonyy9 (jhonyy9):

@jerey45 above you have said B. - why ?

OpenStudy (anonymous):

ok. but the problem is kind of confusing because the part that says "Meg's age is 2 years less than five times Victor's age." is what I mostly don't understand

OpenStudy (anonymous):

you guys. I need like an actual explanation, not to be rude @jerey45 but I want to understand why the answer would be c. Also how would i solve those equations?

OpenStudy (anonymous):

anybody still there?

jhonyy9 (jhonyy9):

try to solve C. for v - ok. ?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so you want me to solve v?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

I think 2?

jhonyy9 (jhonyy9):

right so than you know that v=2 and you know again that m=v+6 so than m= ?

OpenStudy (anonymous):

let me solve it really quick.

OpenStudy (anonymous):

m=8?

OpenStudy (anonymous):

yup

jhonyy9 (jhonyy9):

yes sure so than v=2 and m=8

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

thank you so much guys. I figured out how to medal and i will medal you guys

jhonyy9 (jhonyy9):

@SenpaiMax do you understand it now ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i will fan and medal :D

OpenStudy (anonymous):

I fan you @jerey45 i didn't know that you can only do one best response. can you help me with another question?

OpenStudy (anonymous):

yeah i'll do my best

OpenStudy (anonymous):

ok thxs

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A pair of equations is shown below: y = 8x − 9 y = 4x − 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer. (5 points)

OpenStudy (anonymous):

ummm..... let me think one minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I mostly want to know how to solve those equations

jhonyy9 (jhonyy9):

@SenpaiMax so do you know what mean substitution or elimination method ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

May you explain what that is?

jhonyy9 (jhonyy9):

y=8x-9 y=4x-1 - substitution method is when the value of y from first equation so in this case you will substitute in place of y inside second equation and will calculi easy the value of x - ok ?

OpenStudy (anonymous):

it is like replacing something

OpenStudy (anonymous):

ok. I kind of get it

jhonyy9 (jhonyy9):

- elimination method so for eliminate adding same terms of these equations you need multiplie first or second by -1 and assume the same terms so one will be eliminated and solve for second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do I do then?

jhonyy9 (jhonyy9):

what method you wan us here now ?

OpenStudy (anonymous):

maybe substituion?

jhonyy9 (jhonyy9):

ok. so than replace the first value of y in place of y inside second equation and calculi x

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

can you give me an example please?

jhonyy9 (jhonyy9):

so there is y = kkkkk y = pppp so just replace kkkkk in place of y inside second equation - how you make it ?

jhonyy9 (jhonyy9):

so will be kkkkk= pppp yes ?

OpenStudy (anonymous):

oh ok. i understand. so for my question, what would i do. Do i just put 8x - 9 = 4x -1?

OpenStudy (anonymous):

is that it?

jhonyy9 (jhonyy9):

yes - sure - this is easy

OpenStudy (anonymous):

Ok. Thanks. so i how would i show my work. this math is a bit confusing to em

jhonyy9 (jhonyy9):

wait --- so than calculi the value of x

OpenStudy (anonymous):

what do you mean by calculi?

jhonyy9 (jhonyy9):

8x-9 = 4x-1 8x-4x= -1+9 can you continue ?

OpenStudy (anonymous):

how did you do that?

jhonyy9 (jhonyy9):

(8-4)x = 9-1

OpenStudy (anonymous):

so it would be 4x = 8?

jhonyy9 (jhonyy9):

yes and divide both sides by 4 and what will get for x= ?

OpenStudy (anonymous):

so pretty much what you did was group the variable x with the other one? and x=2?

jhonyy9 (jhonyy9):

yes sure x=2 is right so now you need calculi y=?

OpenStudy (anonymous):

how would you do that?

OpenStudy (anonymous):

i know how to solve for the x now, but y?

jhonyy9 (jhonyy9):

y=4x-1 but you know that x=2 so than y= 4*2 -1 = ?

OpenStudy (anonymous):

oh ok. y =7?

jhonyy9 (jhonyy9):

yes and finish

OpenStudy (anonymous):

ok. thanks. what about part b?

jhonyy9 (jhonyy9):

do you understand it now how solve same system ?

OpenStudy (anonymous):

i understand this now and about substitute and elimination

OpenStudy (anonymous):

i get the graphing now. (x,y) would be (2,7)?

OpenStudy (anonymous):

right?

jhonyy9 (jhonyy9):

yes - right ,this is the answer to part b.

OpenStudy (anonymous):

ok. thanks so much @jhonyy9

jhonyy9 (jhonyy9):

np was my pleasure good luck

OpenStudy (anonymous):

thxs. have a nice day/night